# TPT-Bench: A Large-Scale, Long-Term and Robot-Egocentric Dataset for Benchmarking Target Person Tracking

### v1

@@ V-A Platform, Annotation and Labels
Input: A video sequence V, Initial target bounding box $B_{\_}{\text{init}}$; object detector Det; feature extractor Ext, target feature queue Queue, insertion confidence threshold $\tau$
@@
4             $feat=\texttt{Ext}(I_{\_}k,B_{\_}{\text{init}})$
@@ 5             Queue $\leftarrow feat$
6             Return $\{B_{\_}{\text{init}},1.0\}$
@@
8             $\mathcal{D}_{\_}k=$Det($f_{\_}k$)
9             $\mathcal{T}_{\_}k\leftarrow\emptyset$
10             for $d$ in $\mathcal{D}_{\_}k$ do
11                   $feat=\texttt{Ext}(I_{\_}k,d)$
12                   $sim=\frac{\sum_{\_}{i=1}^{N}feat\cdot\texttt{Queue}[i]}{N}$
13                   $\mathcal{T}_{\_}k\leftarrow\mathcal{T}_{\_}k\cup\{d,feat,sim\}$
@@
15            $\{d_{\_}{\text{max}},feat_{\_}{\text{max}},sim_{\_}{\text{max}}\}\leftarrow\max(\mathcal{T}_{\_}k[sim])$
16             if $sim_{\_}{\text{max}}>\tau$ then
17                   Queue $\leftarrow feat_{\_}{\text{max}}$
@@
20       Return $\{d_{\_}{\text{max}},sim_{\_}{\text{max}}\}$
@@
Input: A video sequence V, Initial target bounding box $B_{\_}{\text{init}}$; person tracker Track; feature extractor Ext, target feature $feat_{\_}{\text{tar}}$, damped version of the exponential moving average DEMA, refer to CARPE-ID [5].
@@
1 for $frame$ $I_{\_}k$ in V do
2       $\{d,id\}_{\_}k=\texttt{Track}(I_{\_}k)$
3       $feats=\texttt{Ext}(I_{\_}k,\{d\}_{\_}k)$
@@
6             $id_{\_}{\text{tar}}=\arg\max\texttt{IoU}(B_{\_}{\text{init}},\{d\}_{\_}k)$
7             $feat_{\_}{\text{tar}}=feats[id_{\_}{\text{tar}}]$
@@
9            Return $\{B_{\_}{\text{init}},1.0\}$
@@
13                   $sim_{\_}{\text{max}}=\texttt{CalCosSim}(feat_{\_}{\text{tar}},feats[id_{\_}{\text{tar}}])$
14                   $feat_{\_}{\text{tar}},\lambda_{\_}s=\texttt{DEMA}(feat_{\_}{\text{tar}},feats[id_{\_}{\text{tar}}])$
15                   Return $\{\{d\}_{\_}k[id_{\_}{\text{tar}}],sim_{\_}{\text{max}}\}$
@@
19                   $sims=\texttt{CalCosSim}(feats,feat_{\_}{\text{tar}})$
20                   $id_{\_}{\text{max}}=\arg\max sims$
21                   $sim_{\_}{\text{max}}=sims[id_{\_}{\text{max}}],d_{\_}{\text{max}}=\{d\}_{\_}k[id_{\_}{\text{max}}]$
22                   if $sim_{\_}{\text{max}}>\lambda_{\_}s$ then
23                         $id_{\_}{\text{tar}}=id_{\_}{\text{max}}$
@@
25                   Return $\{d_{\_}{\text{max}},sim_{\_}{\text{max}}\}$
@@
Input: A video sequence V, Initial target bounding box $B_{\_}{\text{init}}$; person tracker Track; feature extractor Ext, target classifier Cls, feature queue $\mathbb{S}$, ID switch threshold $\tau_{\_}{\text{sw}}$, ReID threshold $\tau_{\_}\text{re}$.
@@
1 for $frame$ $I_{\_}k$ in V do
2       $\{d,id\}_{\_}k=\texttt{Track}(I_{\_}k)$
3       $feats=\texttt{Ext}(I_{\_}k,\{d\}_{\_}k)$
@@
6             $id_{\_}{\text{tar}}=\arg\max\texttt{IoU}(B_{\_}{\text{init}},\{d\}_{\_}k)$
@@
10             $\mathbb{S}\leftarrow\{feats[id_{\_}{\text{tar}}],1\}$
11             for $i$ in $\{id\}_{\_}k\setminus{id_{\_}{\text{tar}}}$ do
@@ 14            Sample $m$ from $\mathbb{S}$, train Cls with $m$
15             $sim_{\_}{\text{max}}=\texttt{Cls}(feats[id_{\_}{\text{tar}}]),d_{\_}{\text{max}}=\{d\}_{\_}k[id_{\_}{\text{tar}}]$
16             if $sim_{\_}{\text{max}}<\tau_{\_}{\text{sw}}$ then
17                   $id_{\_}{\text{tar}}=-1$
@@
19             Return $\{d_{\_}{\text{max}},sim_{\_}{\text{max}}\}$
@@ 23             $sims=\texttt{Cls}(feats)$
24             $id_{\_}{\text{max}}=\arg\max sims$
25             $sim_{\_}{\text{max}}=sims[id_{\_}{\text{max}}],d_{\_}{\text{max}}=\{d\}_{\_}k[id_{\_}{\text{max}}]$
26             if $sim_{\_}{\text{max}}>\tau_{\_}\text{re}$ then
27                   $id_{\_}{\text{tar}}=id_{\_}{\text{max}}$
@@
29             Return $\{d_{\_}{\text{max}},sim_{\_}{\text{max}}\}$
@@
Method AO F-Score AMR FPS
ByteTrack [46]+Target-ReID
CARPE-ID w/ R18 22.04 21.35 0.76 21.2
CARPE-ID [5] w/ KPR 53.61 52.65 11.65 12.0
RPF-ReID w/ R18 50.19 48.38 5.18 20.3
RPF-ReID [4] w/ KPR 65.31 63.26 16.16 11.5
RPF-ReID+OCL [4] w/ R18 57.67 55.71 11.29 18.5
RPF-ReID+OCL [4] w/ Parts-R18 65.11 64.45 8.13 7.1
TABLE IV: Performance comparison of different target-ReID methods on the TPT-Bench. The best result is indicated in bold, while the second-best result is underscored.
Input: A video sequence V, Initial target bounding box $B_{\_}{\text{init}}$; person tracker Track; feature extractor Ext, target classifier Cls, feature queue $\mathbb{S}$, replay buffer $\mathbb{L}$, ID switch threshold $\tau_{\_}{\text{sw}}$, ReID threshold $\tau_{\_}\text{re}$.
@@
1 for $frame$ $I_{\_}k$ in V do
2       $\{d,id\}_{\_}k=\texttt{Track}(I_{\_}k)$
3       $feats=\texttt{Ext}(I_{\_}k,\{d\}_{\_}k)$
4       $patches=\texttt{crop}(I_{\_}k,\{d\}_{\_}k)$
@@
6             $id_{\_}{\text{tar}}=\arg\max\texttt{IoU}(B_{\_}{\text{init}},\{d\}_{\_}k)$
@@
10             $\mathbb{S}\leftarrow\{feats[id_{\_}{\text{tar}}],1\}$, $\mathbb{L}\leftarrow\{patches[id_{\_}{\text{tar}}],1\}$
11             for $i$ in $\{id\}_{\_}k\setminus{id_{\_}{\text{tar}}}$ do
@@ 16             Sample $l$ from $\mathbb{L}$, train Ext with $l$
17             $sim_{\_}{\text{max}}=\texttt{Cls}(feats[id_{\_}{\text{tar}}]),d_{\_}{\text{max}}=\{d\}_{\_}k[id_{\_}{\text{tar}}]$
18             if $sim_{\_}{\text{max}}<\tau_{\_}{\text{sw}}$ then
19                   $id_{\_}{\text{tar}}=-1$
@@
21             Return $\{d_{\_}{\text{max}},sim_{\_}{\text{max}}\}$
@@ 25             $sims=\texttt{Cls}(feats)$
26             $id_{\_}{\text{max}}=\arg\max sims$
27             $sim_{\_}{\text{max}}=sims[id_{\_}{\text{max}}],d_{\_}{\text{max}}=\{d\}_{\_}k[id_{\_}{\text{max}}]$
28             if $sim_{\_}{\text{max}}>\tau_{\_}{\text{re}}$ then
29                   $id_{\_}{\text{tar}}=id_{\_}{\text{max}}$
@@
31             Return $\{d_{\_}{\text{max}},sim_{\_}{\text{max}}\}$
@@
V-B Method Flows of Evaluated TPT Algorithms
In this paper, we evaluate three types of trackers that are capable of performing TPT, including SOT, MOT+Target-ReID, and Detection+ReID. Given only the initial location of the target person’s bounding box, SOT trackers [35, 36, 37, 38, 39, 40, 41, 42, 44, 45] infer estimated bounding box and its confidence of sequent frames with deep neural networks. Due to the large number of SOT methods evaluated in this paper, providing pseudocode for each approach is infeasible within the scope of this work. The pipeline of “Detection+ReID” is shown in Algorithm 1, where a queue is employed to store high-confidence target features and identify the target person by selecting the detected candidate (by YOLOX [49]) with the highest feature similarity.
For the type of “MOT+Target-ReID,” we evaluate different target-ReID methods based on the MOT results from ByteTrack [46]. In this paper, we evaluate “CAPRE-ID [5],” “RPF-ReID [4],” and “RPF-ReID+OCL [4].” The pipeline of “CAPRE-ID” is illustrated in Algorithm 2. In this approach, the target feature is treated as a distribution, and a damped exponential moving average method is employed to update this distribution. This technique helps update the target appearance model gradually, preventing overfitting to newly observed, similar target features.
RPF-ReID [4] treats the target appearance model as a classifier, training it with features collected online. Ridge regression is employed for classification. A simplified pipeline is presented in Algorithm 3. In addition to classifier training, they introduce the fine-tuning of the feature extractor using a replay buffer. This buffer stores online-collected image patches and corresponding labels (target and non-target). A key design element is the consolidation mechanism, driven by replay-based OCL techniques. This mechanism enhances the diversity of stored samples. In the evaluation, we employ the Reservoir [50] method for consolidation. The resulting approach is called “RPF-ReID+OCL,” with a simplified diagram shown in Algorithm 4.
Note that when using part-based ReID features, such as KPR [47], which can infer part-based features and part visibility, we construct multiple part-based target appearance models for target confidence estimation. The final confidence is then obtained by averaging all individual estimations. All methods are evaluated on a PC with an Intel® Core™ i9-12900K CPU and NVIDIA GeForce RTX 3090.
V-C RPF-ReID+OCL with Human Joints
According to [4], an enhanced approach involves leveraging human joint information to explicitly represent partially visible human bodies. Using estimated visibilities, they transform local features from the feature extractor (R18) into features associated with specific body parts. Additionally, part-based classifiers are developed for target confidence estimation. The results, referred to as ”RPF-ReID+OCL w/ parts-R18,” are presented in Table IV, alongside results from other target-ReID methods for comparison.
We observe a significant improvement in F-Score by +8.74, although the AMR decreases by -3.16 compared to ”RPF-ReID+OCL w/ R18.” Furthermore, the FPS drops to 7.1 due to additional computation of human joint estimation and the fine-tuning of the part-based feature extractor.
V-D Why Pushing A Cart for Data Collection
Over the last decade, several datasets have been proposed to enhance robots’ ability to navigate in human environments [17, 29, 30]. In addition to JRDB [17] that highlights perception challenges, some datasets [29, 30] aim to teach robots to navigate socially in human-populated environments. For example, the SCAND dataset [29] includes sequential robot sensor data collected through teleoperated demonstrations to learn socially aware and independent navigation behaviors in human-populated settings. While teleoperated demonstrations can convey human intentions, they suffer from delays in navigation decision-making and are not convenient for non-robotics researchers. This is particularly critical in the person-following scenarios, where we need to respond quickly, considering both the target person and the surrounding environment.
In addition to teleoperated demonstrations, the MuSoHu dataset [30] collects data using a sensor-equipped helmet worn by a naturally walking human. However, this method overlooks the inherent kinematic constraints of robots and the nuances of realistic robot-centric observation, which typically involve increased occlusion and obstruction scenarios encountered in real-world deployments. While the aforementioned data collection methodologies effectively capture egocentric observations and the corresponding human decision-making processes—thereby aiding the development of social navigation techniques—we propose an alternative approach. Specifically, our TPT dataset is collected by having a human push a sensor-equipped cart while following, accompanying, or occasionally searching for a leading person. This method not only reflects natural person-following behavior but also considers the actual kinematic constraints of robotic platforms.
V-E More Dataset Examples
We present four example sequences from the dataset, as shown in Figure 5, which capture multi-modal data visualizations, including panoramic images, point clouds, and odometry. While this paper primarily addresses the long-term visual tracking challenges in the TPT task, we believe the dataset can support other important research areas. For instance, multi-modal TPT can leverage point cloud data to overcome the limitations of visual sensors in low-light conditions. Furthermore, since our data collection method reflects human intentions without delay and all collected data is robot-egocentric, it provides an opportunity to develop learning-based policies based on robot-egocentric observations and actions derived from odometry.
V-F Ethical Statement
This dataset was collected by continuously recording a consenting individual moving through public spaces such as cafeterias and shopping malls. Explicit informed consent was obtained from the primary subject before data collection. We acknowledge that other individuals may have been incidentally captured in the footage while appearing in the background in these public areas. Given that the data was collected in publicly accessible spaces where individuals generally do not have a reasonable expectation of privacy, and no personally identifiable information (PII) of bystanders was intentionally recorded or used, we believe that the collection adheres to ethical guidelines. At the time of submission, we have obtained formal Institutional Review Board (IRB) approval.

### v2

@@ 6 Conclusion
Input: A video sequence V, Initial target bounding box $B_{\text{init}}$; object detector Det; feature extractor Ext, target feature queue Queue, insertion confidence threshold $\tau$
@@
4             $feat=\texttt{Ext}(I_{k},B_{\text{init}})$
@@ 5             Queue $\leftarrow feat$
6             Return $\{B_{\text{init}},1.0\}$
@@
8             $\mathcal{D}_{k}=$Det($f_{k}$)
9             $\mathcal{T}_{k}\leftarrow\emptyset$
10             for $d$ in $\mathcal{D}_{k}$ do
11                   $feat=\texttt{Ext}(I_{k},d)$
12                   $sim=\frac{\sum_{i=1}^{N}feat\cdot\texttt{Queue}[i]}{N}$
13                   $\mathcal{T}_{k}\leftarrow\mathcal{T}_{k}\cup\{d,feat,sim\}$
@@
15            $\{d_{\text{max}},feat_{\text{max}},sim_{\text{max}}\}\leftarrow\max(\mathcal{T}_{k}[sim])$
16             if $sim_{\text{max}}>\tau$ then
17                   Queue $\leftarrow feat_{\text{max}}$
@@
20       Return $\{d_{\text{max}},sim_{\text{max}}\}$
@@
Input: A video sequence V, Initial target bounding box $B_{\text{init}}$; person tracker Track; feature extractor Ext, target feature $feat_{\text{tar}}$, damped version of the exponential moving average DEMA, refer to CARPE-ID (Rollo et al., 2024).
@@
1 for $frame$ $I_{k}$ in V do
2       $\{d,id\}_{k}=\texttt{Track}(I_{k})$
3       $feats=\texttt{Ext}(I_{k},\{d\}_{k})$
@@
6             $id_{\text{tar}}=\arg\max\texttt{IoU}(B_{\text{init}},\{d\}_{k})$
7             $feat_{\text{tar}}=feats[id_{\text{tar}}]$
@@
9            Return $\{B_{\text{init}},1.0\}$
@@
13                   $sim_{\text{max}}=\texttt{CalCosSim}(feat_{\text{tar}},feats[id_{\text{tar}}])$
14                   $feat_{\text{tar}},\lambda_{s}=\texttt{DEMA}(feat_{\text{tar}},feats[id_{\text{tar}}])$
15                   Return $\{\{d\}_{k}[id_{\text{tar}}],sim_{\text{max}}\}$
@@
19                   $sims=\texttt{CalCosSim}(feats,feat_{\text{tar}})$
20                   $id_{\text{max}}=\arg\max sims$
21                   $sim_{\text{max}}=sims[id_{\text{max}}],d_{\text{max}}=\{d\}_{k}[id_{\text{max}}]$
22                   if $sim_{\text{max}}>\lambda_{s}$ then
23                         $id_{\text{tar}}=id_{\text{max}}$
@@
25                   Return $\{d_{\text{max}},sim_{\text{max}}\}$
@@
RPF-ReID (Ye et al., 2024) treats the target appearance model as a classifier, training it with features collected online. Ridge regression is employed for classification. A simplified pipeline is presented in Algorithm 3. In addition to classifier training, they introduce the fine-tuning of the feature extractor using a replay buffer. This buffer stores online-collected image patches and corresponding labels (target and non-target). A key design element is the consolidation mechanism, driven by replay-based OCL techniques. This mechanism enhances the diversity of stored samples. In the evaluation, we employ the Reservoir (Chaudhry et al., 2019) method for consolidation. The resulting approach is called “RPF-ReID+OCL,” with a simplified diagram shown in Algorithm 4.
Note that when using part-based ReID features, such as KPR (Somers et al., 2025), which can infer part-based features and part visibility, we construct multiple part-based target appearance models for target confidence estimation. The final confidence is then obtained by averaging all individual estimations. All methods are evaluated on a PC with an Intel® Core™ i9-12900K CPU and NVIDIA GeForce RTX 3090.
Table 4: Performance comparison of different target-ReID methods on the TPT-Bench. The best result is indicated in bold, while the second-best result is underscored.
Method AO F-Score AMR FPS
ByteTrack (Zhang et al., 2022) + Target-ReID
CARPE-ID w/ R18 21.46 20.72 0.62 21.2
CARPE-ID w/ KPR (Rollo et al., 2024) 56.41 55.28 9.84 12.0
RPF-ReID w/ R18 48.38 46.87 3.89 20.3
RPF-ReID w/ KPR (Ye et al., 2024) 67.44 66.06 16.26 11.5
RPF-ReID+OCL w/ R18 (Ye et al., 2024) 60.25 58.58 11.33 18.5
RPF-ReID+OCL w/ Parts-ResNet18 (Ye et al., 2024) 68.11 68.54 8.15 7.09
Table 5: Detailed annotated scenarios, merged categories.
Category Scenarios
Retail & Shopping bakery, book store, bookstore, clothing store, distribution center, gift store, retail store mobile phone store, supermarket, shopping mall, shopping mall walkway, pickup point, photo booth
Food & Dining student canteen, canteen, food court, coffee bar, street food market, restaurant, fast-food restaurant, Chinese restaurant, McDonald’s, KFC, tea shop
Corridors & Hallways research institute hallway, school hallway, campus hallway, indoor corridor, indoor lobby, connected corridor, classroom corridor, campus corridor, corridor, hallway, breezeway, colonnade, lobby
Roadways & Walkways campus roadway, campus, covered walkway, footbridge, sidewalk, pedestrian walkway, pedestrian crosswalk, urban walkway, street crossing, crosswalk, zebra crosswalk
Vertical Transport elevator, escalator
Outdoor Squares & Plazas square, outdoor plaza, campus courtyard
Educational Rooms classroom, lecture hall
Transport Hubs bus stop, subway station, metro
Entertainment & Leisure game center, game store, pool hall, amusement center
Laboratory & Office robotics laboratory, meeting room, office
Appendix B: RPF-ReID+OCL with human joints
According to Ye et al. (2024), an enhanced approach involves leveraging human joint information to explicitly represent partially visible human bodies. Using estimated visibilities, they transform local features from the feature extractor (R18) into features associated with specific body parts. Additionally, part-based classifiers are developed for target confidence estimation. The results, referred to as ”RPF-ReID+OCL w/ parts-R18,” are presented in Table 4, alongside results from other target-ReID methods for comparison.
We observe a significant improvement in F-Score by +7.86, although the AMR decreases by -3.18 compared to ”RPF-ReID+OCL w/ R18.” Furthermore, the FPS drops to 7.1 due to additional computation of human joint estimation and the fine-tuning of the part-based feature extractor.
Input: A video sequence V, Initial target bounding box $B_{\text{init}}$; person tracker Track; feature extractor Ext, target classifier Cls, feature queue $\mathbb{S}$, ID switch threshold $\tau_{\text{sw}}$, ReID threshold $\tau_{\text{re}}$.
@@
1 for $frame$ $I_{k}$ in V do
2       $\{d,id\}_{k}=\texttt{Track}(I_{k})$
3       $feats=\texttt{Ext}(I_{k},\{d\}_{k})$
@@
6             $id_{\text{tar}}=\arg\max\texttt{IoU}(B_{\text{init}},\{d\}_{k})$
@@
10             $\mathbb{S}\leftarrow\{feats[id_{\text{tar}}],1\}$
11             for $i$ in $\{id\}_{k}\setminus{id_{\text{tar}}}$ do
@@ 14            Sample $m$ from $\mathbb{S}$, train Cls with $m$
15             $sim_{\text{max}}=\texttt{Cls}(feats[id_{\text{tar}}]),d_{\text{max}}=\{d\}_{k}[id_{\text{tar}}]$
16             if $sim_{\text{max}}<\tau_{\text{sw}}$ then
17                   $id_{\text{tar}}=-1$
@@
19             Return $\{d_{\text{max}},sim_{\text{max}}\}$
@@ 23             $sims=\texttt{Cls}(feats)$
24             $id_{\text{max}}=\arg\max sims$
25             $sim_{\text{max}}=sims[id_{\text{max}}],d_{\text{max}}=\{d\}_{k}[id_{\text{max}}]$
26             if $sim_{\text{max}}>\tau_{\text{re}}$ then
27                   $id_{\text{tar}}=id_{\text{max}}$
@@
29             Return $\{d_{\text{max}},sim_{\text{max}}\}$
@@
Input: A video sequence V, Initial target bounding box $B_{\text{init}}$; person tracker Track; feature extractor Ext, target classifier Cls, feature queue $\mathbb{S}$, replay buffer $\mathbb{L}$, ID switch threshold $\tau_{\text{sw}}$, ReID threshold $\tau_{\text{re}}$.
@@
1 for $frame$ $I_{k}$ in V do
2       $\{d,id\}_{k}=\texttt{Track}(I_{k})$
3       $feats=\texttt{Ext}(I_{k},\{d\}_{k})$
4       $patches=\texttt{crop}(I_{k},\{d\}_{k})$
@@
6             $id_{\text{tar}}=\arg\max\texttt{IoU}(B_{\text{init}},\{d\}_{k})$
@@
10             $\mathbb{S}\leftarrow\{feats[id_{\text{tar}}],1\}$, $\mathbb{L}\leftarrow\{patches[id_{\text{tar}}],1\}$
11             for $i$ in $\{id\}_{k}\setminus{id_{\text{tar}}}$ do
@@ 16             Sample $l$ from $\mathbb{L}$, train Ext with $l$
17             $sim_{\text{max}}=\texttt{Cls}(feats[id_{\text{tar}}]),d_{\text{max}}=\{d\}_{k}[id_{\text{tar}}]$
18             if $sim_{\text{max}}<\tau_{\text{sw}}$ then
19                   $id_{\text{tar}}=-1$
@@
21             Return $\{d_{\text{max}},sim_{\text{max}}\}$
@@ 25             $sims=\texttt{Cls}(feats)$
26             $id_{\text{max}}=\arg\max sims$
27             $sim_{\text{max}}=sims[id_{\text{max}}],d_{\text{max}}=\{d\}_{k}[id_{\text{max}}]$
28             if $sim_{\text{max}}>\tau_{\text{re}}$ then
29                   $id_{\text{tar}}=id_{\text{max}}$
@@
31             Return $\{d_{\text{max}},sim_{\text{max}}\}$
@@
Appendix C: Annotated scenarios
Ten high-level categories are distilled from the full set of labelled scenarios listed, which are shown in Table 5.